# Honeybee: Decentralized Peer Sampling with Verifiable Random Walks for Blockchain Data Sharding

### v1

@@ 6. Related Work
6.1. p2p Networks
@@
Whether the purpose is for enhancing blockchain data availability sampling or not, researchers have made extensive effort to improve the security and efficiency of p2p networks. In Castro et al. (Castro et al., 2002), the authors proposed secure routing against attacks that target p2p network message deliveries. The secure routing is achieved with restricted node ID assignment, constrained routing table, and diverse delivery routes. In Baumgart and Mies (Baumgart and Mies, 2007), the authors presented a secure Kademlia key-based routing protocol. They limited free node ID generation with a supervised signature or a crypto puzzle signature and introduced a reliable sibling broadcast. In Coretti et al. (Coretti et al., 2022), the authors designed a byzantine-resilient gossip protocol under the proof-of-stake setting. Under the protocol, nodes build a connected backbone of high-staking nodes. However, light nodes (or nodes with small stakes) are easily sacrificed in this protocol, and stake distribution is not always readily available in a completely decentralized environment. Recent researches inspired by the multi-armed bandit problem (Mao et al., 2020; Wei and Bojja Venkatakrishnan, 2023; Zhang and Bojja Venkatakrishnan, 2023) attempted to reduce the communication latency in p2p networks without losing security by learning from the performance of neighboring peers and update peer selections based on the knowledge. In Kiffer et al. (Kiffer et al., 2021), the authors studied the connectivity and block propagation mechanism of Ethereum’s p2p network. In Król et al. (Król et al., 2023), the authors conducted an analysis on p2p networking requirements for data availability sampling in Ethereum.
6.2. Data Availability Sampling
@@
Plenty of efforts have been made in blockchain layer-2 scaling (Ethereum.org, 2024b, e; Sivaraman et al., 2020; Zhang and Venkatakrishnan, 2023). As a critical component of layer-2 solutions, data availability sampling itself has many important research problems (Neu, 2022). In Hall-Andersen et al. (Hall-Andersen et al., 2023), the authors initiated a cryptographic study of data availability sampling and demonstrated its relation with erasure codes. In Al-Bassam et al. (Al-Bassam et al., 2021), the authors developed a complete fraud and data availability proof scheme in which they claimed that light nodes can have a security guarantee close to the level of a full node under certain assumptions. In Yu et al. (Yu et al., 2020), the authors created coded Merkle trees to improve the protection of light nodes against data availability attacks. In Cao et al. (Cao et al., 2020), the authors presented a decentralized collaborative light-node-only verification mechanism in which light nodes can conduct block verification without the help of a full node. In addition to the advancement in security, there are some researches that focus on lowering the communication and computation overhead on nodes in the data availability setting (Sheng et al., 2021; Kiayias et al., 2020; Bünz et al., 2020; Kiayias et al., 2016).

### v4

@@ 5. Related Work
We have covered some related work on Honeybee in §1. In this section, we will discuss additional related work.
@@
Researchers have made extensive effort to improve the security and efficiency of p2p networks. Protocols like Castro et al. (Castro et al., 2002), S/Kademlia (Baumgart and Mies, 2007), SecureCyclon (Antonov and Voulgaris, 2023), and RAPTEE (Pigaglio et al., 2022) enhance the security of their underlying DHT and Gossip networks by restricting free node ID generation through a centralized authority, cryptographic puzzles, or trusted execution environments. In the Generals’ Scuttlebutt (Coretti et al., 2022), Coretti and colleagues designed a Byzantine-resilient gossip protocol under the proof-of-stake setting. Under the protocol, nodes build a connected backbone of high-staking nodes. However, light nodes (or nodes with small stakes) are easily sacrificed in this protocol, and stake distribution is not always readily available in a completely decentralized environment. In Brahms (Bortnikov et al., 2008), Bortnikov and colleagues presented a Byzantine resistant sampling mechanism in the gossip model. The mechanism, however, relies on the assumption that the ability of malicious nodes to send messages is throttled. In AnonBoot (Matzutt et al., 2020), the authors presented a peer configuration protocol that relies on a proof-of-work chain and requires nodes to advertise themselves on-chain using cryptographic puzzles. Traditional P2P system proposals, such as Tarzan (Freedman and Morris, 2002) and MorphMix (Rennhard and Plattner, 2004), handled attacks based on the assumption that attackers have limited resources (e.g., IP addresses) and that only contiguous ranges of IP addresses are under malicious control. Wang and Borisov proposed Octopus (Wang and Borisov, 2012), a secure lookup strategy in p2p networks, and the strategy relies on a centralized authority to remove malicious nodes from networks. Panchenko and colleagues proposed NISAN (Panchenko et al., 2009), a protocol that conducts anonymous lookups for Chord DHTs using aggregated greedy search and search value hiding. The protocol, however, is vulnerable to information leak attacks and does not offer sufficient protection against Sybil attacks (Wang et al., 2010; Matzutt et al., 2020). Recent researches inspired by the multi-armed bandit problem (Zhang and Bojja Venkatakrishnan, 2023a; Mao et al., 2020) attempted to reduce the communication latency in p2p networks without losing security by learning from the performance of neighboring peers and update peer selections based on the knowledge. In Kiffer et al. (Kiffer et al., 2021), the authors studied the connectivity and block propagation mechanism of Ethereum’s p2p network. In Vedula et al. (Vedula et al., 2023), the authors formalized the p2p topology construction in Ethereum as a game between miners. In Król et al. (Król et al., 2023), the authors conducted an analysis on p2p networking requirements for data availability sampling in Ethereum.
Random walks in p2p networks have been well-discussed by researchers since the 2000s. In contrast to the reason for Honeybee employing verifiable random walks, a plethora of works on p2p random walks primarily focused on resource searching, load balancing, and topology formation for p2p networks (Bisnik and Abouzeid, 2005; Gkantsidis et al., 2004; Zhong et al., 2008; Navaz and Fiaz, 2015; Kwong and Tsang, 2005). In ShadowWalker (Mittal and Borisov, 2009), nodes achieve anonymous communication through random walks with neighborhood verification from shadow nodes, but it relies on the assumption that Sybil nodes constitute at most 20% of the network and possess limited capabilities. In Bisnik et al. (Bisnik and Abouzeid, 2007), the authors improved resource searching efficiency with adaptive random walks that take advantage of the feedback from previous searches. In Massoulié et al. (Massoulié et al., 2006), the authors approached the problem of estimating the number of peers in a p2p network with random walks. In Das Sarma et al. (Sarma et al., 2015), the authors attempted to perform node sampling with minimized round complexity and message complexity using continuous random walks under limited adversarial behaviors. In the Dandelion Bitcoin network protocol (Bojja Venkatakrishnan et al., 2017), the authors applied random walks before message broadcast to protect the anonymity of the transaction parties. Augustine et al. (Augustine et al., 2022) use multiple random walks to achieve peer sampling for DHTs by introducing a rather complicated process of building and dismantling hypercubes of nodes. However, their random walks can be susceptible to tampering and only allow a maximum of $\frac{n}{polylog(n)}$ Byzantine nodes in a network of $n$ nodes.
@@
Honeybee can be applied to enhance blockchain data availability sampling at the network layer through its node sampling capability. Plenty of efforts have been made in blockchain layer-2 scaling (Ethereum.org, 2024a, c; Sivaraman et al., 2020; Zhang and Bojja Venkatakrishnan, 2023b; Ma et al., 2025). As a critical component of layer-2 solutions, data availability sampling itself has many important research problems (Neu, 2022). In Hall-Andersen et al. (Hall-Andersen et al., 2023), the authors initiated a cryptographic study of data availability sampling and demonstrated its relation with erasure codes. In Al-Bassam et al. (Al-Bassam et al., 2021), the authors developed a complete fraud and data availability proof scheme in which they claimed that light nodes can have a security guarantee close to the level of a full node under certain assumptions. In Yu et al. (Yu et al., 2020), the authors created coded Merkle trees to improve the protection of light nodes against data availability attacks. In Cao et al. (Cao et al., 2020), the authors presented a decentralized collaborative light-node-only verification mechanism in which light nodes can conduct block verification without the help of a full node. In addition to the advancement in security, there are some researches that focus on lowering the communication and computation overhead on nodes in the data availability setting (Sheng et al., 2021; Kiayias et al., 2020; Bünz et al., 2020; Kiayias et al., 2016).